OpenStudy (anonymous):

Find the missing factor. Write your answer in exponential form. \[1=?*3^5\]

OpenStudy (agent0smith):

Just solve for the ? like you would for x, divide both sides by 3^5\[\Large \frac{ 1 }{3^5 }=?\]

OpenStudy (usarmy3947):

did this help? ^

OpenStudy (kittiwitti1):

Let's make the ? be x.\[1=a\times 3^5\]\[\frac{1}{3^5}=a\]\[a=0.0041152263\]\[a\approx0.004\]I don't know if you need units but add them if you do.

OpenStudy (kittiwitti1):

I mean make the ? be a, sorry. You could also write the answer another way:\[1=a\times3^5\]\[\frac{1}{3^5}=a\]\[a=3^{-5}\]